# Measuring Agents in Production

### v1

@@ Abstract
AI agents are already operating in production across many industries, yet there is limited public understanding of the technical strategies that make deployments successful. We present the first large-scale systematic study of AI agents in production, surveying 306 practitioners and conducting 20 in-depth case studies via interviews across 26 domains. We investigate why organizations build agents, how they build them, how they evaluate them, and what the top development challenges are. We find that production agents are typically built using simple, controllable approaches: 68% execute at most 10 steps before requiring human intervention, 70% rely on prompting off-the-shelf models instead of weight tuning, and 74% depend primarily on human evaluation. Reliability remains the top development challenge, driven by difficulties in ensuring and evaluating agent correctness. Despite these challenges, simple yet effective methods already enable agents to deliver impact across diverse industries. Our study documents the current state of practice and bridges the gap between research and deployment by providing researchers visibility into production challenges while offering practitioners patterns emerging from successful deployments.
@@ 1 Introduction
Large language models (LLMs) have enabled a new class of software systems: AI Agents. We define AI agents as systems that combine foundation models with optional tools, memory, and reasoning capabilities to autonomously execute multi-step tasks [Weng_LLM_Powered_Autonomous_Agents_2023, OpenAI_Docs_Agents, GoogleCloud_WhatAreAIAgents, IBMThink_WhatAreAIAgents, NVIDIA_Intro_LLM_Agents_2023]. Recent research has demonstrated exciting potential for AI agents across domains from drug discovery [Huang2025_biomni, Swanson2025VirtualLab, liu2025drugagentautomatingaiaideddrug], algorithmic innovation [novikov2025alphaevolvecodingagentscientific, barbariansgateaiupending, shojaee2025llmsrscientificequationdiscovery] to general AI scientist [lu2024aiscientistfullyautomated, gottweis2025aicoscientist, yamada2025aiscientistv2workshoplevelautomated]. This excitement extends beyond research prototypes to production systems. Companies report agents actively running in diverse domains far beyond the initial coding focus [CursorAgent2024, ClaudeCode2025, OpenAICodex, Windsurf], including finance [IACPM_McKinsey_GenAI_Credit_2025, Verhagen2025AgenticAIFinancialCrime, Capgemini2025AIAgentsFraud, Deloitte_AgenticAI_Banking_2025], insurance [Allianz2025ClaimQueueAgenticAI, EconomistImpact2025UnderwritingFutureAIInsurance, KPMG_Intelligent_Insurance_2025, Everest_Group_AgenticAI_Insurance_2025], and education [Linsenmayer2025LeveragingAIForSEN, Kakar2024JillWatsonScaling, MichiganRoss2025GoogleAgenticAIVirtualTA].
However, despite widespread excitement about agent potential [Wang_2024, Maslej2025AIIndexReport, CapgeminiCRI2025RiseAgenticAI_Web, PwC_AIAgentSurvey_2025], many question their real value and success in production [xue2025illusionprogressassessingcurrent, Reuters2025GartnerAgenticAICanceled, cmu2025usability, MITMediaLab_NANDA_2025]: whether agents can deliver on their promise and where their future lies. For example, a recent study reports that 95% of agent deployments fail [MITMediaLab_NANDA_2025]. This stark contrast between the promise of agents and their high failure rate raises fundamental questions about what separates successful deployments from failures.
Unfortunately, little information is publicly available on how production agents are built. Why do some agents succeed while others fail? What requirements must agents meet for production deployment? We lack a systematic understanding of what methods enable successful agents in the real world. Researchers have little visibility into real-world constraints and practices. Are agents failing because models are not capable enough, or are there other factors at play? Without understanding how agents actually work in production, researchers risk addressing the wrong problems. Meanwhile, practitioners lack consensus on how to build reliable agents and would benefit from understanding how the industry approaches these fast-evolving systems.
@@
To address this knowledge gap, we present Measuring Agents in Production (MAP), the first large-scale systematic study of AI agents in production. We study the practices of developers and teams behind successful real-world systems via four research questions ( s):
@@
What are the applications, users, and requirements of agents?
@@
What models, architectures, and techniques are used to build deployed agents?
@@
How are agents evaluated for deployment?
4.
What are the top challenges in building deployed agents?
We answer these questions through an online survey with 306 responses and 20 in-depth interviews with agent development teams, capturing technical details of production agent development. Our survey respondents are practitioners actively building AI agents (Figure 4(a)) across 26 domains, from finance and healthcare to legal (Figure 2). We filter the survey data to 86 responses that explicitly reported systems in production or pilot phases,***Production systems are fully deployed and used by target end users in live operational environments. Pilot systems are deployed to controlled user groups for evaluation, phased rollout, or safety testing. Definitions for other stages appear in Appendix E.2 N5. which we denote as deployed agents, for the results in the main paper. The full dataset with all 306 responses appears in Appendix A. For in-depth interviews, we study 20 cases from teams building deployed agents with real users, spanning organizations from global enterprises to startups (Figure 3). Our case studies and survey capture deployed agent systems serving user bases ranging from hundreds to millions of daily users (Figure 4(c)), providing visibility into high-impact deployments. This dual approach provides the first empirically grounded view of the technical methods, architectural patterns, and operational practices behind AI agents in production.
To provide first-hand data on working practices and real-world constraints, our study focuses on practitioners actively building agents already deployed in production that serve users and deliver measurable value (Figure 1). This naturally biases toward successful deployments and experienced practitioners. Our data capture what works today, providing insight into how the field is developing.
Our study reveals key findings for each research question:
1 : Productivity gains drive agent adoption. We find that 73% of practitioners deploy agents primarily to increase efficiency and decrease time spent on manual tasks (Figure 1). Most systems (93%) serve human users rather than other agents or systems (Figure 5(a)). Regarding requirements, teams prioritize output quality over real-time responsiveness: 66% allow response times of minutes or longer, compared to 34% requiring sub-minute latency (Figure 5(b)). In-depth interviews reveal that organizations tolerate this latency because agents that take minutes to execute still outperform human baselines.
2 : Simple methods and workflows dominate. 70% of interview cases use off-the-shelf models without weight tuning (Figure 6), relying instead on prompting. Teams predominantly select the most capable, expensive frontier models available, as the cost and latency remain favorable compared to human baselines. We find that 79% of surveyed deployed agents rely heavily on manual prompt construction (Figure 7), and production prompts can exceed 10,000 tokens. Production agents favor well-scoped, static workflows: 68% execute at most ten steps before requiring human intervention, with 47% executing fewer than five steps. Furthermore, 85% of detailed case studies forgo third-party agent frameworks, opting instead to build custom agent application from scratch. Organizations deliberately constrain agent autonomy to maintain reliability.
3 : Human verification remains central to evaluation. The majority of deployed survey agents (74%) rely primarily on human-in-the-loop evaluation, and 52% use LLM-as-a-judge (Figure 10(b)). Notably, every interviewed team utilizing LLM-as-a-judge also employs human verification. Because production tasks are highly domain-specific, public benchmarks are rarely applicable. Consequently, 25% of teams construct custom benchmarks from scratch, while the remaining 75% evaluate their agents without formal benchmarks, relying instead on online-tests such as A/B testing or direct expert/user feedback. This pattern reflects the difficulty of automated evaluation for bespoke production tasks.
4 : Reliability is an unsolved challenge. Practitioners focus most on ensuring agent reliability, spanning correctness, latency, and security. The evaluation challenges described in 3 directly impact the ability to verify correctness to achieve reliable deployments. Latency impacts only a small subset (15%) of applications as a deployment blocker, and security represents a manageable concern that most deployed agents mitigate through action and environment constraints.
Our findings reveal that reliability concerns drive practitioners toward simple yet effective solutions with high controllability, including restricted environments, limited autonomy, and human oversight. For example, developers choose human-in-the-loop evaluation over fully automated techniques and manual prompt engineering over automated methods because these approaches offer control, transparency and trustworthiness. Practitioners deliberately trade-off additional agent capability for production reliability, and this design pattern already enables a broad spectrum of applications delivering real-world value (Figure 1). Production agents represent an emerging engineering discipline. Our study documents the current state of practice and bridges the gap between research and deployment: we provide researchers visibility into real-world constraints and opportunities while offering practitioners proven patterns from successful deployments across industries.
@@ 2 Related Work
To our knowledge, we offer the first technical characterization of how agents running in production are built.
Commercial Agent Surveys.
Several prior efforts examine AI (agent) adoption in production from adjacent perspectives. MITMediaLab_NANDA_2025 and mlq2025stateofai study economic viability and executive views on return-on-investment from companies attempting to integrate agents. cmu2025usability analyze marketing materials for 102 commercial agents and conduct user studies with 31 participants, revealing gaps between promised and realized capabilities. Industry reports [PwC_AIAgentSurvey_2025, Capgemini_AI_Agents_2025, PagerDuty_AgenticAI_Survey_2025, McKinsey_SeizingAgenticAI_2025, Microsoft_WTI_2025_FrontierFirm] focus on organizational readiness and market trends. LangChain_State_of_AI_Agents_2024 surveyed over 1,300 professionals on agent motivations, characteristics, and challenges.
| Business Operations |
| Insurance claims workflow automation |
| Customer care internal operations assistance |
| Human resources information retrieval and task assistance |
| Communication Tech, Multi-lingual Multi-dialect |
| Communication automation services |
| Automotive communication services |
| Scientific Discovery |
| Biomedical sciences workflow automation |
| Materials safety and regulatory analysis automation |
| Chemical data interactive exploration |
| Software & Business Operations |
| Data analysis and visualization |
| Enterprise cloud engineer and business assistance |
| Site reliability incident diagnoses and resolution |
| Software products technical question answering |
| Software DevOps |
| Spark version code and runtime migration |
| Software development life cycle assistance end-to-end |
| Software engineer/developer slack support |
| SQL optimization |
| Code auto-completion and syntax error correction |
Our work differs in two key ways: (i) Scope: we study agents actively operating in production; and (ii) Focus: we collect engineering-level technical data from practitioners. Our study complements these prior works and offers a new perspective into successful agents, providing insights into where the field is heading.
Research Agent Literature Survey. Many prior work examine LLM-powered agents from an academic perspective [krishnan2025aiagentsevolution, liu2025advanceschallengesfoundationagents, wang2024survey, luo2025largelanguagemodelagent, PICCIALLI2025128404, plaat2025agenticlargelanguagemodels], providing valuable taxonomies of agent designs and tracing the evolution of key techniques. Other agent surveys focus on specific aspects: evaluation methodologies [yehudai2025surveyevaluationllmbasedagents, Mohammadi_2025], security concerns [agent_security_survey], and multi-agent systems [guo2024largelanguagemodelbased]. In contrast, our work is an empirical study that collects primary data directly from practitioners building and operating deployed agents. We do not synthesize published research via literature reviews; instead, we conduct original survey data collection and in-depth case study interviews to document production practices.
Single-System Studies. Companies publish papers and technical blogs on agentic systems [gottweis2025coscientist, shen2023hugginggptsolvingaitasks, chennabasappa2025llamafirewallopensourceguardrail, prabhakar2025enterprisedeepresearchsteerable, jha2025itbenchevaluatingaiagents, Anthropic_Developing_Computer_Use_2024_misc, Cursor2025TabRL, Cognition2025SWE15FastAgentModel, Anthropic2025MultiAgentResearchSystem] and open-source agent implementations [Block2025Goose, OpenAI2025Codex, GeminiCLI_Docs, Cline2025Homepage, OpenHands2025Homepage], offering glimpses into production-related agents. Each paper naturally focuses on a single system or domain, creating an information gap: we lack understanding of common patterns, shared challenges, and design principles across deployments. We address this gap by surveying professionals across diverse industries to identify recurring engineering practices, architectural decisions, and challenges.
3 Methodology
To understand how production AI agents are built and deployed, we conducted two complementary studies: a large-scale public survey and 20 in-depth case study interviews. We detail our survey design and distribution in Section 3.1, describe our interview protocol and case study selection in Section 3.2, and outline our data processing pipeline in Section 3.3.
3.1 Public Online Survey
We designed and conducted a survey for practitioners building AI agents, with a strong focus on the technical details of their systems. The survey covers system architecture, evaluation, deployment, and operational challenges through 47 questions (listed in full in Appendix E). To capture AI Agents as practitioners understand them, we invited respondents to describe systems they refer to as AI Agents or Agentic AI without imposing a prescriptive definition.
To ensure relevance, we implemented the survey with dynamic branching logic in Qualtrics, where certain questions only appear based on previous responses. For example, the question about the realized benefits of using agents is only shown to respondents who confirmed they chose agentic over non-agentic solutions (O3.1.1 and O3.1 in Appendix Additional Questions). This adaptive structure, combined with the optional nature of all questions, means each question receives a different number of responses. We report the specific sample size (N) for each result throughout Sections 4–7 in figure captions. We asked participants to focus their responses on a single system if they contributed to more than one agent system (Acknowledgment E.1). Full questionnaire details are available in Appendix E, including the branching logic in Figures 26 and 27.
We distributed the survey through multiple channels to reach practitioners across the AI agent ecosystem: the Berkeley RDI Agentic AI Summit [BerkeleyRDI_AgenticAISummit_2025_misc], the AI Alliance Agents-in-Production Meetup [AIAllianceMeetup_5Aug2025_misc], the UC Berkeley Sky Retreat [SkyComputingRetreatsCamps], and professional networks including LinkedIn, Discord, and X. We release the survey on July 28, 2025, with data collection for this paper ending October 29, 2025.
We received 306 valid responses. 294 participants indicated that they have directly contributed to the building and designing of at least one agent system. Our survey respondents are predominantly technical professionals such as software and machine learning engineers (Figure 4(a)). Among 111 respondents who reported their system’s deployment stage, 82% indicated their systems are in production or pilot phases (Figure 4(b)), demonstrating rapid transition from experimental prototypes to real-world deployments.
To maintain focus on production systems, we filtered the data to 86 responses that explicitly reported their systems as being in production or pilot phases. Production systems are fully deployed and used by target end users in live operational environments, while pilot systems are deployed to controlled user groups for evaluation, phased rollout, or safety testing. This filtering excludes development prototypes, research artifacts, and retired systems, as well as participants who did not report their deployment stage. The definitions for all deployment stages appear in Appendix E.2 N5. We denote production and pilot systems as deployed agents. All survey statistics presented in this paper refer to deployed agents unless otherwise stated. Complete data for 306 valid responses across all deployment stages appears in Appendix A.
3.2 In-depth Case Studies
To add qualitative depth to our survey findings, we conducted 20 in-depth case studies through interviews with teams building deployed agents across a diverse range of organizations.
We carefully selected our 20 cases to achieve representative samples across application diversity, organizational maturity, and global reach. We only interviewed systems with real-world users: 14 cases are in full production and 6 cases are in final pilot phases. The cases span five key sectors: business operations (3 cases), software development and operations (7 cases), integrated business and software operations (5 cases), scientific discovery (3 cases), and communication services (2 cases). Notably, these deployments extend well beyond commonly known coding agents or general chatbots, demonstrating the breadth of production agent applications. The systems target both internal employees (5 cases) and external enterprise consumers (15 cases). Our selection includes organizations across maturity levels, from seed-stage startups to large enterprises with global footprints (Figure 3). For companies with multiple agent deployments, we presented only distinct use cases to maximize application diversity. The anonymized case studies and their application domains appear in Table 1, with additional details in Appendix D.
Each interview lasted 30 to 90 minutes and was conducted by teams of 2 to 5 organizationally neutral interviewers. We followed a semi-structured interview protocol covering 11 topic areas (detailed in Appendix D.1), including system architecture, evaluation mechanisms, deployment challenges, operational requirements, and measurable agent values. We anonymized and recorded interviews based on participant preferences, with human note-takers capturing insights. To ensure accuracy, we cross-validated final summaries among all interviewers. Per our confidentiality agreements, we anonymized all data and present findings in aggregate.
3.3 Data Processing
Most survey questions use structured formats (single-select, multi-select, or numeric), requiring minimal post-processing. For the free-text domain keyword responses used in Figure 2, we used LOTUS [lotus], a state-of-the-art unstructured data processing tool, to identify common domain categories and perform classification. This allowed us to normalize phrases from survey responses to representative label sets. For instance, we grouped responses like “healthcare,” “medical,” and “patient monitoring” under a unified “healthcare” category. Details of this normalization process appear in Appendix B.1. All other figures present results from structured survey questions or interview data without requiring automated processing.
As described in Section 3.1, our survey data in the main paper are filtered to deployed (production and pilot) agents, and our interviews specifically select teams building deployed agents. All results presented in Sections 4–7 refer to deployed agents from either survey responses or interviews, which we explicitly denote throughout the paper. Refer to Appendix A for unfiltered full data.
For all figures that include error bars, we report 95% confidence intervals computed using 1,000 bootstrap samples with replacement.
4 1 Results: What Are The Applications, Users, and Requirements of Agents?
We now present findings from our survey and case study interviews across four central research questions. We start by examining motivations for agent adoption (Section 4.1), which agent applications successfully reach deployment (Section 4.2), who uses these systems (Section 4.3), and what latency requirements shape their design (Section 4.4). Understanding these patterns reveals where agentic systems deliver practical value and how they transform real-world applications.
4.1 Motivations for Choosing Agents
Among practitioners with deployed agents who evaluated alternatives, 82.6% prefer the agentic solution for deployment. We define non-agentic alternatives as existing software systems, traditional approaches, or human execution. Figure 1 details the specific benefits reported by these practitioners. The top three drivers all target the reduction of human effort: increasing productivity and efficiency (72.7%), reducing human hours (63.6%), and automating routine labor (50.0%). Conversely, qualitative benefits rank lower. Increasing user satisfaction ranks in the middle tier (37.9%), followed by reducing required human expertise and training (34.8%) and enabling novel technology (33.3%). Accelerating failure response times (18.2%), reducing interdisciplinary knowledge requirements (18.2%), and risk mitigation (12.1%) rank last.
These priorities reflect pragmatic deployment realities. Organizations adopt agents to solve immediate operational problems, such as expert-expensive manual work and insufficient staffing capacity. Productivity gains are straightforward to quantify through human-hour reductions, whereas safety improvements and risk mitigation are harder to verify. This focus aligns with our findings on internal deployments (Section 4.3), where organizations tolerate higher error rates by maintaining human oversight to catch these errors.
The top reported reasons for using agents reveal a trend where certain objectives are more verifiable and measurable. For example, time to complete a task (productivity) is concrete and quantifiable, while risk mitigation benefits takes longer to verify. We examine this measurement gap and its implications for deployed agents in Section 7.1.

### v2

@@ Abstract
LLM-based agents already operate in production across many industries, yet we lack an understanding of what technical methods make deployments successful. We present the first systematic study of Measuring Agents in Production, MAP, using first-hand data from agent developers. We conducted 20 case studies via in-depth interviews and surveyed 306 practitioners across 26 domains. We investigate why organizations build agents, how they build them, how they evaluate them, and their top development challenges. Our study finds that production agents are built using simple, controllable approaches: 68% execute at most 10 steps before human intervention, 70% rely on prompting off-the-shelf models instead of weight tuning, and 74% depend primarily on human evaluation. Reliability (consistent correct behavior over time) remains the top development challenge, which practitioners currently address through systems-level design. MAP documents the current state of production agents, providing the research community with visibility into deployment realities and underexplored research avenues.
@@ 1 Introduction
Large language models enable a new class of software systems—agents—that combine foundation models with tools, memory, and reasoning to autonomously execute multi-step tasks (Wang et al., 2024c). LLM-based agents have gained substantial research interest, demonstrating potential in areas such as drug discovery and scientific discovery (Huang et al., 2025; Novikov et al., 2025; Lu et al., 2024). Industry is also now deploying agents in domains central to society: finance, healthcare, and education (IACPM, McKinsey, 2025; UHS, Inc., 2025; Linsenmayer, 2025).
Despite widespread excitement about agents, studies show agent deployments often fail or underdeliver (Xue et al., 2025; Reuters Staff, 2025; Shome et al., 2025). The stark contrast between the potential of agents and their failures raises the fundamental question of what enables successful agent deployment. The field can only advance collectively through shared understanding of real-world challenges and lessons learned. Yet, unfortunately, little information is publicly available on how production agents are built.
@@
To address this knowledge gap, we present MAP, the first large-scale systematic study of AI agents in production. We study the practices of developers and teams behind successful real-world systems via four research questions (RQs):
RQ1.
What are the applications of agents?
RQ2.
What models, architectures, and methods are used?
RQ3.
How are agents evaluated?
RQ4.
What are the top challenges in deployment?
We answer these questions by conducting 20 in-depth interviews with deployment teams ranging from major technology companies with AI research labs to agent startups (Figure 15), and surveying 306 practitioners actively building agents across 26 domains (Figure 2). We conducted the study from April to November of 2025. We filtered survey responses to 86 systems in production or pilot phases (Figure 14(b)), serving hundreds to millions of daily users (Figure 14(c)). We refer to these production and pilot systems as deployed agents and focus our analysis on them in the main paper. Full unfiltered survey data appear in §D. Our key findings for each RQ are the following:
RQ1: Productivity gains drive agent adoption. Practitioners deploy agents primarily to increase productivity (80%), mainly serving human users rather than other systems. Deployed agents operate in latency-tolerant applications: 66% allow response times of minutes or longer, as agents outperform human baselines even with these latencies.
RQ2: Production agents favor simplicity and control. Teams use off-the-shelf models rather than weight tuning (70%), rely on manual prompt construction (79%), and use static workflows (68% execute 10 steps before human intervention). Organizations deliberately trade capability for controllability to maintain reliability.
RQ3: Human-in-the-loop evaluation dominates. Deployed agents rely primarily on human-in-the-loop evaluation (74%); other methods (e.g., LLM-as-a-judge) serve as complementary verification. Due to limited available benchmarks and challenges in creating them, 75% of teams forgo formal benchmarking, relying instead on A/B testing or expert feedback to improve reliability.
RQ4: Reliability is the top development challenge. Other critical challenges include evaluation (limited by benchmark scarcity and delayed feedback) and security (mitigated via operational constraints).
Our findings suggest an underlying principle: practitioners achieve reliability through best practices in system-level design rather than model-level or algorithmic advances. For example, despite the popularity of RL in research and its benchmark gains, practitioners default to prompting closed-source models because this approach is more robust to model upgrades and more sample-efficient. Thus, teams deliberately choose simple, controllable methods not from lack of sophistication, but because they offer reliable agent performance and fast development cycles.
MAP documents real-world agent practices with first-hand data from practitioners. By sharing deployment data, which are typically kept proprietary, we connect research advances with real-world constraints and challenges. We hope these data-driven insights can help inspire the community to address the underexplored research directions and technical challenges revealed by our study, advancing the field of agents together to deliver value in the real-world. The contributions of this paper are as follows:
@@
First large-scale empirical study of production agents: We conduct 20 in-depth interviews with deployment teams and survey 306 practitioners, including 86 in deployments, providing systematic data on agents serving real users.
@@
Characterization across 17 design dimensions: We provide quantitative and qualitative data on technical decisions, including models, architectures, prompting, evaluation, applications, and operational constraints.
@@
Data-driven insights: We show practitioners achieve reliability through system-level design rather than algorithmic advances. Our data provides foundational evidence to support diverse research directions in agent systems.
@@ 2 Related Work
To our knowledge, we offer the first technical characterization of how agents running in production are built. Additional related work in each category appears in §F.
Commercial Agent Surveys. Prior work examines agent adoption from adjacent perspectives. Challapally et al. (2025) study economic viability from business executives’ views at companies attempting agent integration. Shome et al. (2025) analyze marketing materials and conduct user studies, revealing capability gaps. LangChain (2024) surveyed 1,300 professionals on motivations and challenges. And various consulting reports focus on organizational readiness (PwC, 2025; Sukharevsky et al., 2025). Our work differs in scope (production systems) and focus (engineering-level technical data from practitioners).
Research Literature Surveys. Academic surveys synthesize published research, providing overviews of agent architectures and taxonomies (Wang et al., 2024b; Liu et al., 2025), with specialized surveys on evaluation (Yehudai et al., 2025) and security (He et al., 2025). In contrast, our empirical study collects primary data directly from practitioners building deployed agents through original interviews and surveys rather than synthesizing existing literature.
Single-System Studies. Companies publish on individual systems (Gottweis et al., 2025; Anthropic, 2024; Jackson et al., 2025) and open-source implementations (OpenAI, 2025a; OpenHands, 2025), providing valuable insights into specific systems. We report common patterns across diverse deployments rather than documenting individual systems.
3 Study Methodology
To understand the technical methods enabling the production of agents, we collected data through 20 in-depth interviews and surveyed 306 practitioners, conducting the study from April to November of 2025. However, such a study presents distinct challenges: (1) successful agent deployments with real-world users are scarce, and (2) production agents are typically proprietary, limiting company participation.
We address these challenges through systematic sampling and rigorous study design. Throughout the study, we prioritize diversity (domains and organization sizes) and transparency (documenting selection criteria and limitations).
3.1 Data Collection #1: In-depth Case Studies
We conducted 20 case studies through interviews with ML and software engineers.
Participants. Rather than selecting all participants upfront, we iteratively expanded our sample guided by emerging findings following snowball sampling (Naderifar et al., 2017). We began with agent teams from major technology companies with significant AI research within our professional network, then systematically broadened across application diversity, organizational maturity, and global reach. This yielded cases spanning five sectors (Table 3), organization sizes from startups to global enterprises (Figure 15), and deployment scales from hundreds to millions of users across multiple continents. All cases represent verified production (14) or pilot (6) deployments serving real users. We anonymized all case studies and refer to each as C01, C02, etc. in result sections (Table 3 in §B).
Semi-structured interviews. We conducted 30-90 minute interviews with teams of 2-5 researchers with no affiliation to the participating organizations following a protocol covering system architecture, evaluation, and deployment challenges (§C.3). The semi-structured format allowed us to explore unexpected themes while maintaining consistency, and each interview is documented under the same protocol.
Analysis. After each interview, researchers documented observed patterns for theme extraction in addition to interview notes. We then applied grounded theory’s (Glaser and Strauss, 1967) open coding across all interviews to identify recurring themes, followed by focused coding to organize themes into broader categories.
3.2 Data Collection #2: Public Online Survey
Survey design. We developed a 47-question survey to confirm our qualitative observations from our case studies and to understand agents deployment at scale and address our four research questions. We piloted the survey with agent development teams outside our study, iteratively refining questions based on feedback. We implemented the final survey with dynamic branching in Qualtrics, where subsequent questions adapt based on prior responses, allowing relevant deep-dives while maintaining completion rates. All questions were optional to respect participants’ freedom to answer. Complete survey questions appear in §G.
Distribution and collection. We distributed the validated survey through (Anonymous) agent-focused conferences, agent online courses, and professional networks (LinkedIn, Discord, X) from July 28 to October 29, 2025, targeting practitioners who build agent systems. We received 306 valid responses from practitioners across diverse technical roles (Figure 14(a)) and 26 application domains (Figure 2).
Data filtering and processing. We filtered responses to 86 data points explicitly in production or pilot phases. All results in §4–7 refer to these as deployed agents; full unfiltered data appear in §D. Most survey questions use structured formats (single-select, multi-select, numeric). For domain classification (Figure 2), the only free-text question, we applied LOTUS (Patel et al., 2025) for semantic aggregation to derive candidate domain categories and employed three independent human annotators for labeling (Cohen’s =0.636); details appear in §B.1.2. For data on categorical comparisons, we report 95% confidence intervals computed using 1,000 bootstrap samples with replacement.
3.3 Study Limitations
Despite best efforts to obtain broad sample coverage, our study has inherent limitations in scope and sampling. Geographically, case study teams are concentrated in the Americas, with a few in Europe. Participation bias affects both data sources: agent teams accepted interview invitations based on company policies, while survey respondents likely skew toward our professional networks despite multi-channel distribution. Multi‑month data collection broadened system coverage but may also have introduced temporal bias, which we mitigated in part by emphasizing themes that remained consistent over time. Thus, we view MAP as an initiative towards documenting the leading edge of production agents practice for open research rather than a comprehensive coverage of all agents development globally.
4 RQ1 Results: What Are the Applications of Agents?
We present findings from our study on why organizations build agents, which applications reach deployment, who uses them, and what requirements shape their design.
4.1 Motivation for Building Agents
Deployed agents primarily target measurable productivity gains. Among surveyed practitioners with deployed agents, 80% cite increased productivity, and 72% cite reduced human task-hours (Figure 1). Benefits that are harder to quantify emerge less frequently in current deployments, such as risk mitigation (12%) and reducing the need for interdisciplinary expertise (18%). Productivity-focused applications offer straightforward success metrics: interviewed teams commonly measure productivity gains by comparing total time to completion between agents and alternative systems. In contrast, operational improvements like risk mitigation require extended validation periods before benefits become measurable. Among practitioners who evaluated alternative systems for the same objectives, 83% prefer agents over non-agentic solutions (software or human execution).